OpenStudy (anonymous):

Solve 12x^2 + 14x - 6 = 0

OpenStudy (anonymous):

Have you learned how to use the quadratic formula yet? It's useful when solving when it's harder to factor (or not factorable).

OpenStudy (anonymous):

The answer is x = -3/2 and -1/3

OpenStudy (anonymous):

I lied :P my answer is incorrect lol thats what i get for rushing lol

OpenStudy (anonymous):

From your question it looks like synthetic division can be used o.o

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

hero (hero):

12x^2 + 14x - 6 = 0 2(6x^2 + 7x - 3) = 0 6x^2 + 7x - 3 = 0 After finding two numbers that add to get 7 but multiply to get (6)(-3) or -18 6x^2 + 9x - 2x - 3 = 0 3x(2x + 3) -1(2x + 3) = 0 (2x + 3)(3x - 1) = 0 x = -3/2, 1/3